### v1

@@ Abstract
MDS matrices play a critical role in the design of diffusion layers for block ciphers and hash functions due to their optimal branch number. Involutory and orthogonal MDS matrices offer additional benefits by allowing identical or nearly identical circuitry for both encryption and decryption, leading to equivalent implementation costs for both processes. These properties have been further generalized through the notions of semi-involutory and semi-orthogonal matrices. While much of the existing literature focuses on identifying efficiently implementable MDS candidates or proposing new constructions for MDS matrices of various orders, this work takes a different direction. Rather than introducing novel constructions or prioritizing implementation efficiency, we investigate structural relationships between the generalized variants and their conventional counterparts. Specifically, we establish nontrivial interconnections between semi-involutory and involutory matrices, as well as between semi-orthogonal and orthogonal matrices. Exploiting these relationships, we show that the number of semi-involutory MDS matrices can be directly derived from the number of involutory MDS matrices, and vice versa. A similar correspondence holds for semi-orthogonal and orthogonal MDS matrices. We also examine the intersection of these classes and show that the number of $3\times 3$ MDS matrices that are both semi-involutory and semi-orthogonal coincides with the number of semi-involutory MDS matrices over $\mathbb{F}_{2^{m}}$. Furthermore, we derive the general structure of orthogonal matrices of arbitrary order $n$ over $\mathbb{F}_{2^{m}}$. Based on this generic form, we provide a closed-form expression for enumerating all $3\times 3$ orthogonal MDS matrices over $\mathbb{F}_{2^{m}}$. Finally, leveraging the aforementioned interconnections, we present explicit formulae for counting $3\times 3$ semi-involutory MDS matrices and semi-orthogonal MDS matrices.
@@ 1 Introduction
    We revisit the problem of enumerating $3\times 3$ semi-involutory MDS matrices over $\mathbb{F}_{2^{m}}$. By leveraging the interconnection with involutory MDS matrices and applying known enumeration results from [12], we derive the enumeration formula for $3\times 3$ semi-involutory MDS matrices over $\mathbb{F}_{2^{m}}$ through a more direct and concise approach. Similarly, we derive the number of $3\times 3$ semi-orthogonal MDS matrices over $\mathbb{F}_{2^{m}}$.
@@
    We extend our enumeration to $4\times 4$ matrices by computing the number of semi-involutory MDS matrices over $\mathbb{F}_{2^{m}}$ for $m=3,4,\ldots,8$, thereby extending the previously known results for $m=3,4$ [16] to larger finite field. Finally, Tables 4 and 5 present a comparison of the counts of $3\times 3$ and $4\times 4$ MDS matrices with various structural properties.
@@ 3.3 The Interconnection between Semi-Involutory and Semi-Orthogonal Matrices
$\alpha_{1}=\alpha_{2}=\ldots\alpha_{n},\text{ and }\beta_{1}=\beta_{2}=\ldots\beta_{n}$.
@@
$M_{1}=\begin{pmatrix}1&1&1&1\\ 1&\alpha^{5}&\alpha&\alpha^{4}\\ 1&\alpha^{4}&\alpha^{10}&\alpha^{2}\\ 1&\alpha^{8}&\alpha^{5}&\alpha^{11}\end{pmatrix}$
@@
From Tables 1 and 3, we can observe that the number of $4\times 4$ MDS matrices that are both semi-involutory and semi-orthogonal is equal to the number of $4\times 4$ semi-involutory MDS matrices over $\mathbb{F}_{2^{3}}$. This equivalence arises because there are no non-symmetric semi-involutory MDS matrices of order $4$ over $\mathbb{F}_{2^{3}}$ of the form $M_{1}$. In contrast, among the semi-involutory MDS matrices of the form $M_{1}$ of order $4$ over $\mathbb{F}_{2^{4}}$, we find $60768$ non-symmetric matrices and $11088$ symmetric matrices. This discrepancy leads to a different count for the matrices over the field $\mathbb{F}_{2^{4}}$.
@@ 4 On the Structure of Orthogonal MDS Matrices
$T_{1}=\left\{0,m_{12}+1,(m_{12}+1)^{-1}m_{12}+1,(m_{12}+1)^{-1}m_{12}+m_{12}\right\}.$
We now demonstrate that the cardinality of $T_{1}$ is $4$ for any $m_{12}\in\mathbb{F}_{2^{m}}\setminus\left\{0,1\right\}$. To establish this, we equate all pairs of elements in $T_{1}$ and solve for $m_{12}$. The resulting cases are as follows:
@@
We now demonstrate that the cardinality of $T_{3}$ is $6$ for any values of $m_{11}\notin\mathbb{F}_{2^{m}}\setminus\{0,1\}$ and $m_{12}\notin\left\{0,1,m_{11},m_{11}+1\right\}$. To establish this, we equate all pairs of elements in $T_{3}$ and solve for $m_{11}$ or $m_{12}$. By analyzing these $15$ cases, we can say that $T_{3}$ contains precisely six elements. Therefore, we have $2^{m}-6$ many choices for $m_{21}$.

### v2

@@ Abstract
MDS matrices play a critical role in the design of diffusion layers for block ciphers and hash functions due to their optimal branch number. Involutory and orthogonal MDS matrices offer additional benefits by allowing identical or nearly identical circuitry for both encryption and decryption, leading to equivalent implementation costs for both processes. These properties have been further generalized through the notions of semi-involutory and semi-orthogonal matrices. While much of the existing literature focuses on identifying efficiently implementable MDS candidates or proposing new constructions for MDS matrices of various orders, this work takes a different direction. Rather than introducing novel constructions or prioritizing implementation efficiency, we investigate structural relationships between the generalized variants and their conventional counterparts. Specifically, we establish nontrivial interconnections between semi-involutory and involutory matrices, as well as between semi-orthogonal and orthogonal matrices. Exploiting these relationships, we show that the number of semi-involutory MDS matrices can be directly derived from the number of involutory MDS matrices, and vice versa. A similar correspondence holds for semi-orthogonal and orthogonal MDS matrices. We also examine the intersection of these classes and show that the number of $3\times 3$ MDS matrices that are both semi-involutory and semi-orthogonal coincides with the number of semi-involutory MDS matrices over $\mathbb{F}_{2^{m}}$. Furthermore, we derive the general structure of orthogonal matrices of arbitrary order $n$ over $\mathbb{F}_{2^{m}}$. Finally, leveraging the aforementioned interconnections, we present an alternative and direct derivation of the explicit formulae for counting $3\times 3$ semi-involutory MDS matrices and $3\times 3$ semi-orthogonal MDS matrices.
@@ 1 Introduction
More specifically, [2] shows that certain Cauchy and Vandermonde based MDS constructions also satisfy the semi-orthogonal property, and it initiates a study of circulant matrices with semi-involutory and semi-orthogonal structure. In [4], the authors further analyze circulant matrices over finite fields and relate the MDS property to the trace of an associated diagonal matrix. In [3], they characterize $3\times 3$ semi-involutory MDS matrices and enumerate them over finite fields of characteristic 2. Independently, [16] derives counting formula for $3\times 3$ semi-involutory MDS matrices over $\mathbb{F}_{2^{m}}$, and additionally provides algorithms to test the semi-involutory and semi-orthogonal properties, detailed structural results for circulant matrices with these properties, and non-existence results for circulant MDS matrices under the same structural constraints.
@@
    We revisit the enumeration of $3\times 3$ semi-involutory and semi-orthogonal MDS matrices over $\mathbb{F}_{2^{m}}$. Exploiting their correspondence with involutory MDS matrices and known enumeration results from [12], we provide a direct derivation of the closed-form count for $3\times 3$ semi-involutory MDS matrices over $\mathbb{F}_{2^{m}}$. In an analogous way, we derive the closed-form count for $3\times 3$ semi-orthogonal MDS matrices over $\mathbb{F}_{2^{m}}$. Together with previously known formulas and those proved in this work, these results yield a consolidated set of explicit enumeration formulas for several algebraically constrained $3\times 3$ MDS classes, summarized in Table 1.
@@
    We extend our enumeration to $4\times 4$ matrices by computing the number of semi-involutory MDS matrices over $\mathbb{F}_{2^{m}}$ for $m=3,4,\ldots,8$, thereby extending the previously known results [16] for $m=3,4$ to larger finite field. Finally, Tables 5 and 6 present a comparison of the counts of $3\times 3$ and $4\times 4$ MDS matrices with various structural properties.
Table 1: Enumeration formulas for $3\times 3$ MDS matrices over $\mathbb{F}_{2^{m}}$
Matrix Type Number of Matrices
$3\times 3$ MDS $(2^{m}-1)^{5}(2^{m}-2)(2^{m}-3)(2^{2m}-9\cdot 2^{m}+21)$ [15]
$3\times 3$ Involutory MDS $(2^{m}-1)^{2}(2^{m}-2)(2^{m}-4)$ [12]
$3\times 3$ Semi-involutory MDS $(2^{m}-1)^{5}(2^{m}-2)(2^{m}-4)$ [3, 16] [This work, Corollary 2]
$3\times 3$ Orthogonal MDS $(2^{m}-2)(2^{m}-3)(2^{m}-4)$ [This work, Theorem 4.2]
$3\times 3$ Semi-orthogonal MDS $(2^{m}-1)^{5}(2^{m}-2)(2^{m}-3)(2^{m}-4)$ [16] [This work, Theorem 4.3]
$3\times 3$ MDS with both Semi-involutory and Semi-orthogonal $(2^{m}-1)^{5}(2^{m}-2)(2^{m}-4)$ [This work, Theorem 3.10]
@@ 3.3 The Interconnection between Semi-Involutory and Semi-Orthogonal Matrices
$\alpha_{1}=\alpha_{2}=\cdots=\alpha_{n},\text{ and }\beta_{1}=\beta_{2}=\cdots=\beta_{n}$.
@@
$M_{1}=(c_{ij})=\begin{pmatrix}1&1&1&1\\ 1&\alpha^{5}&\alpha&\alpha^{4}\\ 1&\alpha^{4}&\alpha^{10}&\alpha^{2}\\ 1&\alpha^{8}&\alpha^{5}&\alpha^{11}\end{pmatrix}$
@@
From Tables 2 and 4, we can observe that the number of $4\times 4$ MDS matrices that are both semi-involutory and semi-orthogonal is equal to the number of $4\times 4$ semi-involutory MDS matrices over $\mathbb{F}_{2^{3}}$. This equivalence arises because there are no non-symmetric semi-involutory MDS matrices of order $4$ over $\mathbb{F}_{2^{3}}$ of the form $M_{1}$. In contrast, among the semi-involutory MDS matrices of the form $M_{1}$ of order $4$ over $\mathbb{F}_{2^{4}}$, we find $60768$ non-symmetric matrices and $11088$ symmetric matrices. This discrepancy leads to a different count for the matrices over the field $\mathbb{F}_{2^{4}}$.
@@ 4 On the Structure of Orthogonal MDS Matrices
[…]
@@
[…]